In the reaction Mg(OH)2 + 2HCl → MgCl2 + 2H2 O, you were given 15



grams of Mg(OH)2 and 25 grams of HCl. Which of the two reactants is the



limiting reagent? How much MgCl2 will be produced in the reaction? (Mg(OH)2



= 58.32 g/mol, HCl = 36.46 g/mol, MgCl2



= 95.32 g/mol)




1
Expert's answer
2022-03-25T15:01:04-0400

Solution:

Calculate moles of each reagent:

Moles of Mg(OH)2 = [15 g Mg(OH)2] × [1 mol Mg(OH)2 / 58.32 g Mg(OH)2] = 0.2572 mol Mg(OH)2

Moles of HCl = [25 g HCl] × [1 mol HCl / 36.46 g HCl] = 0.6857 mol HCl


Balanced chemical equation:

Mg(OH)2 + 2HCl → MgCl2 + 2H2O

According to stoichiometry:

1 mol of Mg(OH)2 reacts with 2 mol of HCl

Thus, 0.2572 mol of Mg(OH)reacts with:

[0.2572 mol Mg(OH)2] × [2 mol HCl / 1 mol Mg(OH)2] = 0.5144 mol HCl

However, initially there is 0.6857 mol of HCl (according to the task)

Therefore, Mg(OH)2 acts as limiting reagent and HCl is excess reagent


According to stoichiometry:

1 mol of Mg(OH)2 produces 1 mol of MgCl2

Thus, 0.2572 mol of Mg(OH)2 produces:

[0.2572 mol Mg(OH)2] × [1 mol MgCl2 / 1 mol Mg(OH)2] = 0.2572 mol MgCl2


Calculate the mass of MgCl2:

The molar mass of MgCl2 is 95.21 g/mol

Therefore,

Mass of MgCl2 = [0.2572 mol MgCl2] × [95.21 g MgCl2 / 1 mol MgCl2] = 24.488 g MgCl2 = 24.5 g MgCl2

Mass of MgCl2 = 24.5 g


Answers:

The limiting reagent is Mg(OH)2

24.5 grams of MgCl2 will be produced in the reaction